OpenStudy (anonymous):

4x^3-17x^2-42x/x^2y-13xy+42y simplify

OpenStudy (jhannybean):

\[\large \frac{4x^3-17x^2-42x}{x^2y-13xy+42y}\]

OpenStudy (anonymous):

yes, now it needs to be simplified =]

OpenStudy (jhannybean):

I just needed to see it written in equation format first lol. One min. :P

OpenStudy (anonymous):

thank you!

jhonyy9 (jhonyy9):

4x^3-17x^2-42x so in case of this what can you factorizing out ?

OpenStudy (jhannybean):

All i've got so far is that.. You can factor out an x from the numerator, and a y from the denominator. \[\large \frac{x(4x^2-17x-42)}{y(x^2-13x+42)}\]

OpenStudy (jhannybean):

Now can you not solve them separately? The numerator and denominator?

OpenStudy (anonymous):

well. i dont know lol. Its mulitple choice but none of it is even close =/ other than the x and y being factored out

OpenStudy (jhannybean):

We an factor out the bottom though :) so that's a plus. What two numbers add to give -13 and multiply to give +42?

OpenStudy (anonymous):

-7 -6

OpenStudy (jhannybean):

Good. \[\large \frac{x(4x^2-17x-42)}{y(x-7)(x-6)}\] So far...

jhonyy9 (jhonyy9):

ok and the numerator ?

OpenStudy (anonymous):

factor out

OpenStudy (jhannybean):

first we can split the 4x^(2) into (x )(4x ) Then analyzing the factors of 42, we understand that +7 and -6 = -42, and adding them together gives you 13, BUT we have a 4 present also, so 13+4 = 17. Knowing this, we an simplify 4x^(2) -17x -42 into \[\large (x-6)(4x+7)\]

OpenStudy (anonymous):

does the x-6 on top and bottom cancel

OpenStudy (jhannybean):

Yep it cancels out.\[\large \frac{x(4x^2-17x-42)}{y(x-7)(x-6)}= \frac{x\cancel{(x-6)}(4x+7)}{y(x-7)\cancel{(x-6)}}\] and what do you get?

OpenStudy (anonymous):

do you bring the 4x outside?

OpenStudy (jhannybean):

Nope.

OpenStudy (anonymous):

nvm lol

OpenStudy (anonymous):

thank you!

OpenStudy (jhannybean):

You can simplify the numerator and denominator by distributing the x and y back into their respective places.

OpenStudy (jhannybean):

and no problem :)

OpenStudy (anonymous):

are you good at restrictions? lol

OpenStudy (jhannybean):

Depends, haha.,

OpenStudy (anonymous):

f(x)= x+1/ x^2+5x-6

OpenStudy (anonymous):

gotta find the restriction

OpenStudy (jhannybean):

How do you simplify x^(2) +5x - 6? What two numbers add together to give 5, but multiply to give -6

OpenStudy (anonymous):

-1 6

OpenStudy (anonymous):

so is that the restriction? -1,6

OpenStudy (jhannybean):

Alright, so you have \[\large \frac{x+1}{(x+6)(x-1)}\] solving for the denominator, set both = 0 \[\large x+6=0\]\[\large x-1=0\] and solving for these x values will give you yourrestriction.

OpenStudy (anonymous):

so x= -6 x=1

OpenStudy (jhannybean):

mmhmm :)

OpenStudy (jhannybean):

and to double check this, plug these values back into your original function \[\large f(x) = \frac{x+1}{x^2+5x-6}\]\[\large f(1)=\frac{2}{0} = \text{undefined}\]\[\large f(-6)= \frac{-6+1}{(-6)^2+5(-6)-6}=\frac{-5}{0}= \text{undefined}\]

OpenStudy (anonymous):

thanks!

OpenStudy (jhannybean):

np :)

OpenStudy (anonymous):

i dont know how you write the problems our on here, can you write this out for me?

OpenStudy (anonymous):

x+2/7-x * x^2-9x+4/x^2-5x+6

OpenStudy (anonymous):

the x+2 is divided by the 7-x

OpenStudy (anonymous):

and same for the other side

OpenStudy (jhannybean):

Use the equation editor :P